Update at 22:52:54(BST): The issue has now been resolved with this server.We are presently investigating a fault affecting the database server on this machine.
13th April 2012, at 10:29PM
13th April 2012, at 9:52PM
Update at 11:28:38(BST): The engineers found a kernel issue and have updated this. After updating, the server rebooted fine and everything is back up and running as expected. We apologise for the inconvenience caused.
Update at 11:00:01(BST): The server is back up and running and the engineers are currently looking at what has caused the issue.
Update at 10:31:27(BST): An engineer is at the data centre and there will be a further update within the next 30 minutes.
Update at 09:40:27(BST): The reboot and diagnostics is taking longer than expected. Engineers are still dealing with the issue and will bring the services back online as soon as possible.
We are currently rebooting Prohost14 due to apache and sshd locking up. We're currently rebooting and looking into the causes of this and the server should be back online shortly.
We apologise for the inconvenience caused.
13th April 2012, at 9:25AM
13th April 2012, at 10:28AM
Update at 13:53:35(BST): Despite the extra memory being placed in the server, we've had various issues with it being recognised/working correctly. We hope to address this ASAP in the future. For now we've got the server back up running stable without the extra memory.
--
The memory chips in server14 have been replaced, this should hopefully aid in ruling out an issue with one of the ones that was present and causing some stability issues.
the disks are currently running full checks before a final reboot on this.
this has taken a bit longer than anticipated due to run disk checks.
10th April 2012, at 12:05PM
10th April 2012, at 12:53PM
Update at 09:53:51(BST): The issue has now been resolved with this server.Websites and Databases are now back online. Apologies for any inconvenience caused.
We are currently scanning the disks on this server to repair any potential issues. Apologies for any inconvenience caused.
10th April 2012, at 9:42AM
10th April 2012, at 8:53AM
Update at 12:05:44(BST): The issue has now been resolved with this server.All services on the max server are now back on line. Thank you for your patience. Apologies for any inconvenience caused.
Update at 11:35:38(BST):
Our engineers are now running a full check on the disks, before the system can be rebooted properly. We thank you for your patience.
Update at 09:12:25(BST):
Our engineers attempted to reboot the server at 8am this morning. Sadly, this did not bring the server back online. On site engineers are now look into the the issue.
Apologies for any inconvenience caused.
This will affect websites and emails on the personal account type on max.34sp.com only.
10th April 2012, at 9:10AM
10th April 2012, at 11:05AM
Update at 08:56:29(BST): The issue has now been resolved with this server.Update at 08:35:29(BST): Phone support will be unavailable until 9:00am today. We apologise for any inconvenience. E-mail support will be available as usual: support@34sp.com.
10th April 2012, at 8:03AM
10th April 2012, at 7:56AM
Update at 14:50:39(BST): At approximately 5.40PM today we noticed a degradation in network performance. This was immediately tracked down and resolved by around 6.00PM.
Users should not have experienced total outages, just slowdowns in accessing services.
Monitoring has shown no re-occurrence of this in the last 7 days and is now resolved.
5th April 2012, at 6:06PM
12th April 2012, at 1:50PM
Update at 09:39:22(BST): Update at 08:25:14(BST): The engineers have restarted the main server itself which has now come back up and all VPS servers contained within are starting up. Please allow 5-10 minutes for all services to start up as normal. If you do notice any issues with your server afterwards please email in to support@34sp.com so we can investigate.
The main VPS server vps23 is currently unreachable which will be affecting all VPS and sites contained within. Our engineers have been notified and are investigating into the matter now.
Please accept our apologies for any inconvenience caused and rest assured we are doing all we can to get this resolved as speedily as possible.
2nd April 2012, at 8:10AM
13th April 2012, at 8:39AM
Update at 15:32:18(BST): The issue has now been resolved with this server.
28th March 2012, at 3:25PM
28th March 2012, at 2:32PM
Update at 15:26:23(BST): This maintenance was completed as planned last night.
27th March 2012, at 9:12PM
28th March 2012, at 2:26PM